If there are five multiple choice questions, each with four possible answers what is the probability of...
1) getting none of them correct?
2) getting all of the questions correct?
3) getting at least one of the questions wrong?
4) getting your first incorrect answer on the fourth question?

Answer by edjones(8007)

1) getting none of them correct?
(3/4)^5 = 243/1024
.
2) getting all of the questions correct?
(1/4)^5 = 1/1024
.
3) getting at least one of the questions wrong?
1 - 1/1024 = 1023/1024
.
4) getting your first incorrect answer on the fourth question?
(1/4)^3 * 3/4 = 3/256
.
Ed